# Bayesian Low-Rank LeArning (Bella): A Practical Approach to Bayesian Neural Networks

### v1

@@ Abstract
Computational complexity of Bayesian learning is impeding its adoption in practical, large-scale tasks. Despite demonstrations of significant merits such as improved robustness and resilience to unseen or out-of-distribution inputs over their non-Bayesian counterparts, their practical use has faded to near insignificance. In this study, we introduce an innovative framework to mitigate the computational burden of Bayesian neural networks (BNNs). Our approach follows the principle of Bayesian techniques based on deep ensembles, but significantly reduces their cost via multiple low-rank perturbations of parameters arising from a pre-trained neural network. Both vanilla version of ensembles as well as more sophisticated schemes such as Bayesian learning with Stein Variational Gradient Descent (SVGD), previously deemed impractical for large models, can be seamlessly implemented within the proposed framework, called Bayesian Low-Rank LeArning (Bella). In a nutshell, i) Bella achieves a dramatic reduction in the number of trainable parameters required to approximate a Bayesian posterior; and ii) it not only maintains, but in some instances, surpasses the performance of conventional Bayesian learning methods and non-Bayesian baselines. Our results with large-scale tasks such as ImageNet, CAMELYON17, DomainNet, VQA with CLIP, LLaVA demonstrate the effectiveness and versatility of Bella in building highly scalable and practical Bayesian deep models for real-world applications.
@@ 1 Introduction
Bayesian deep learning [1] provides mechanisms for building more robust predictive models—i) robustness to adversarial attacks; ii) unseen or out-of-distribution data—and iii) a theoretical framework for estimating model uncertainty [2, 3, 4, 5, 6]. In particular, quantifying uncertainty enables more reliable decision-making and facilitates identifying potential model vulnerabilities in uncertain regions of the input space. Consequently, embracing Bayesian deep neural networks (BNNs) represents a significant stride towards building more reliable and trustworthy AI systems for various real-world applications (e.g., autonomous driving, medical image analysis, etc.). Unfortunately, their practical use is encumbered by their computational complexity.
Unlike traditional alternatives with point estimates—a single set of model parameters mapping inputs to outputs— BNNs learn the distribution of model parameters to offer a distribution over possible predictions. Consider a neural network $f(\mathbf{x},\boldsymbol{\theta})$ with input $\mathbf{x}$ and weights $\boldsymbol{\theta}$ and a prior distribution on network weights $p(\boldsymbol{\theta})$. The likelihood function $p(\mathcal{D}|\boldsymbol{\theta})$ determines how well the network with weights $\boldsymbol{\theta}$ fits the data $\mathcal{D}$ [7]. Bayesian inference integrates the likelihood and the prior using Bayes’ theorem to derive a posterior distribution over the space of weights, given by $p\left(\boldsymbol{\theta}|\mathcal{D}\right)=\frac{p\left(\mathcal{D}|\boldsymbol{\theta}\right)p\left(\boldsymbol{\theta}\right)}{p(\mathcal{D})}$ and seek to compute the predictive distribution
@@
Despite evidence to support the adoption of Bayesian learning, their widespread use is hindered by several challenges; a primary issue is the intractability of the posterior distribution $p(\boldsymbol{\theta}|\mathcal{D})$ in practical learning tasks. The exact solution for the posterior, even for networks of moderate size, is impractical. Because of network complexity and the high-dimensional integral of the resulting denominator. This necessitates the use of approximations.
Figure 1: Train and Test Error (%) landscapes of CAMELYON17. Training landscape demonstrates the learned 3 particle approximation of the modes of the posterior from a pre-trained model $\boldsymbol{\theta}$—modification of parameters in the constrained region ($\Delta\boldsymbol{\theta}$) leads to approaching posterior modes in inference. Interestingly, the test landscape shows a shift in particle performance, reflecting the intrinsic distribution shift in CAMELYON17 test set. Here, we observe a key benefit of our Bella approximation compared to a point estimate—while a single parameter particle (e.g. $\boldsymbol{\theta}+\Delta\boldsymbol{\theta}_{1}$) do not generalize well, Bayesian prediction with Equation 1, effectively an average over multiple parameter settings, leads to better performance.
In this paper, we consider approximations of the posterior as ensemble or Stein Variational Gradient Descent (SVGD) but provide a practical method to learn BNNs for large-scale, learning tasks.
Effectively, SVGD employs a mechanism to separate the parameters of the ensembles (i.e. particles), ensuring the representation of alternative modes in the posterior distribution. But computational demands of the approximation, especially the memory demands for storing all parameters, hinder adoption in large-scale models such as CLIP [8]. However, a recent conjecture [9] suggests that the permutation of optimal model parameters can lead to models that are not functionally different. Further, the path between modes may be linear. Therefore, we consider spawning parameter particles (models) by adding a linear interpolation to a pre-trained model’s parameters. But, it would still require updating a large number of model parameters and we are still confronted by computation complexity. The recent conjecture in [10] suggests modes may yet exist in the constrained region in the parameter space. Therefore, we propose only adding a linear interpolation to a portion of the model parameters and, subsequently, only learning them as illustrated in Figure 1 for the CAMELYON17 task—a benchmark task for out-of-distribution evaluations. We summarise our key contributions as:
@@
  • We propose a new Bayesian learning framework, Bella, for SVGD approximation of the posterior—spawning particles or models, by exploiting the availability of pre-trained models, with linear interpolation of a constrained set of model parameters for learning with SVGD to approximate the posterior.
@@
  • Our approach more efficiently captures the complexity and multi-modality of the solution space compared to current SVGD but at a fraction of the cost—we observe on-par performance in terms of uncertainty estimation, performance improvement, and robustness, updating only 0.3% or less parameters.
@@
  • We demonstrate Bella to perform on par or better than baselines—ensembles or current SVGD implementations. Bella consistently outperforms the non-BNN counterparts on 8 datasets; image classification (ImageNet, CIFAR 10/100), Out-of-distribution (DomainNet, Camleyon17111Notably, our approach sets a new state-of-the-art for the CAMELYON17 (see Table 6 for more details) dataset challenge according to the leaderboard [11].  [12], CIFAR-10-C), and VQA as well as adversarial robustness. Notably, our approach, on a multi-modal architecture such as LlaVa [13], leads to improved performance and uncertainty estimation highly correlated with human confidence.
@@ 2 Related Work
Increasing network complexity has led to advances in parameter-efficient fine tuning and recent advances in exploring these methods for Bayesian learning, predominantly for language models. We discuss these in the context of our work.
@@
Parameter-Efficient Fine Tuning. In contrast to fine-tuning all parameters, recent research proposed inserting adapters in between existing neural layers to reduce the number of trainable parameters and, subsequently, the compute (GPU consumption) [14, 15, 16]. Hu et al. [17] use a bottleneck structure to impose a low-rank constraint on the weight updates, named LoRA. The key functional difference is that LoRA can be merged with the main weights during inference, thus avoiding the introduction of any latency whilst significantly reducing the number of parameters.
@@
Fine-Tuning Approaches and Bayesian Deep Learning.  Previous research investigating the application of fine-tuning approaches for BNNs have predominantly focused on large language models (LLMs), e.g.  [18, 19, 20]. Notably, marking a departure from the conventional methods relying primarily on tuning the network’s parameters, these studies chose to define priors and approximate posterior over low rank attention weights. Concurrently, Yang et al. [20] introduced the concept of Laplace LoRA to incorporate Bayesian concepts to enhance the calibration of fine-tuned LLMs. However, Laplace’s method relies on a Gaussian approximation of the posterior distribution. Whilst this can be effective for unimodal and symmetric distributions, the approach does not fully encapsulate the intricacies of more complex posteriors, particularly in neural networks where the posterior has multimodality and asymmetry [21]. Deep ensembles [22] typically perform better in practice compared with variational and Laplace methods, due to their ability to capture multiple modes. When employing fine-tuning, a direct application of ensembling for LoRAs was considered in [23]. Some interpretations of deep ensembles suggest that they approximate gradient flows in function spaces and that building desirable properties into an ensemble (such as repulsive behavior), is possible [24]. SVGD [25], can be viewed in a similar vein. However, while these more sophisticated, repulsive, ensembling approaches are highly impractical in the pre-training phase, we argue that their expressivity can be brought to bear precisely in tandem with low-rank fine-tuning, which is the viewpoint we adopt in this contribution.
@@
Building upon these foundations, our research represents a pioneering effort to apply the principles of repulsive ensemble-based low-rank fine-tuning to computer vision. In particular, we bridge pre-training and fine-tuning phases with recent conjectures on mode connectivity [9]. Our methodology not only capitalizes on the efficiency of fine-tuning techniques, e.g. [26, 27, 28] but also innovatively addresses the scalability challenges inherent to BNNs. Overall, our work sets a new precedent in applying Bayesian approaches to computer vision tasks by offering a scalable and efficient framework for enhancing model performance.
3 Background on Stein Variational Gradient Descent (SVGD) Approximations for a Bayesian Posterior
Here, we provide: i) a brief overview of methods to address the intractability of the posterior distribution $p(\boldsymbol{\theta}|\mathcal{D})$ in Bayesian learning tasks; and ii) we elaborate on our motivation for seeking a practical method to learn BNNs for large-scale tasks with SVGD.
Bayesian inference techniques have been integral to the development of neural networks, with a rich history underscored by previous works [29, 1, 30, 31, 32]. These methodologies provide a structured framework for generating reliable and interpretable estimates of uncertainty. However, the focus of prior research on Bayesian neural networks has predominantly been on the initial pretraining stages rather than on subsequent fine-tuning processes [33, 34]. This emphasis on pretraining presents significant challenges in terms of scalability and resource requirements, limiting the application of Bayesian neural networks to smaller datasets and network architectures with a minimal number of parameter particles. Consequently, the scalability issues hinder the extension of Bayesian neural networks to larger datasets, such as ImageNet [35], and more complex network architectures, like CLIP [8], which involve millions of parameters.
@@ 3 Background on Stein Variational Gradient Descent (SVGD) Approximations for a Bayesian Posterior
Variational Inference (VI) [36, 32] and Markov Chain Monte Carlo (MCMC) [29, 31] are two primary approximate Bayesian inference frameworks. The former substitutes the true posterior with a tractable alternative while the latter involves sampling. However, accurately computing the posterior with either MCMC or VI becomes computationally infeasible when dealing with large-scale networks containing millions of parameters. Although approximations can be obtained more efficiently with VI [36], VI is also demonstrably too restrictive to resemble the multi-modality of the true posterior and suffers from mode collapse [21].
@@
Stein Variational Gradient Descent (SVGD) [25] is an alternative approximate Bayesian technique which combines the strengths of MCMC and VI by transporting a set of parameter particles to fit the true posterior distribution, while encouraging diversity among the particles, by incorporating a repulsive term in the parameter updates. This diversity prevents the mode collapse and enables learning multiple models to represent various patterns in the data. Using $n$ samples from the posterior (i.e. parameter particles), SVGD modifies the gradient descent as:
$\boldsymbol{\theta}_{i}=\boldsymbol{\theta}_{i}$ $-\epsilon_{i}\hat{\boldsymbol{\phi}}{}^{*}(\boldsymbol{\theta}_{i})\text{\quad with}\quad\hat{\boldsymbol{\phi}}{}^{*}(\boldsymbol{\theta})=\sum_{j=1}^{n}\big{[}k(\boldsymbol{\theta}_{j},\boldsymbol{\theta})\nabla_{\boldsymbol{\theta}_{j}}\log p(\boldsymbol{\theta}_{j}|\mathcal{D})-\frac{\gamma}{n}\nabla_{\boldsymbol{\theta}_{j}}k(\boldsymbol{\theta}_{j},\boldsymbol{\theta})\big{]}\,.$ (2)
@@
4 Bayesian Low-Rank Learning (Bella) for SVGD
The problem with the current SVGD approach in large deep neural networks is its huge computational cost. This renders it infeasible to train efficiently and to scale to a sufficient number of parameter particles for accurately approximating the posterior distribution, which currently remains coarse. In this work, we propose to capitalize on the low-rank representations of fine-tuning in order to construct a practical and scalable variant of SVGD. Consider any dense layer, for which there is a fixed pre-trained weight matrix $\boldsymbol{\theta}_{0}\in\mathbb{R}^{d_{1}\times d_{2}}$ with $d_{1},d_{2}$ the corresponding numbers of hidden units. We consider $n$ low-rank perturbations of $\boldsymbol{\theta}_{0}$ as

### v2

# Bayesian Low-Rank LeArning (Bella): A Practical Approach to Bayesian Deep Learning

@@ Abstract
Computational complexity of Bayesian learning is impeding its adoption in practical, large-scale tasks, despite demonstrations of significant merits such as improved robustness and resilience to unseen or out-of-distribution inputs over their non-Bayesian counterparts. Although, Deep ensemble methods (Seligmann et al. 2024; Lakshminarayanan, Pritzel, and Blundell 2017) have proven to be highly effective for Bayesian deep learning, their practical application is hindered by substantial computational cost. In this study, we introduce an innovative framework to mitigate the computational burden of ensemble Bayesian deep learning. We explore a more feasible alternative, inspired by the recent success of low-rank adapters, we introduce Bayesian Low-Rank LeArning (Bella). We show, i) Bella achieves a dramatic reduction in the number of trainable parameters required to approximate a Bayesian posterior; and ii) it not only maintains, but in some instances, surpasses the performance–in accuracy and out-of-distribution generalisation–of conventional Bayesian learning methods and non-Bayesian baselines. Our extensive empirical evaluation in large-scale tasks such as ImageNet, CAMELYON17, DomainNet, VQA with CLIP, LLaVA demonstrate the effectiveness and versatility of Bella in building highly scalable and practical Bayesian deep models for real-world applications.
@@ 1 Introduction
Figure 1: Train and Test Error (%) landscapes of CAMELYON17. Training landscape demonstrates the learned 3 particle approximation of the modes of the posterior from a pre-trained model $\boldsymbol{\theta}$—modification of parameters in the constrained region ($\Delta\boldsymbol{\theta}$) leads to approaching posterior modes in inference. Interestingly, the test landscape shows a shift in particle performance, reflecting the intrinsic distribution shift. Here, we observe a key benefit of our Bella approximation compared to a point estimate—while a single parameter particle (e.g. $\boldsymbol{\theta}+\Delta\boldsymbol{\theta}_{1}$) do not generalize well, Bayesian prediction with Equation 1, effectively an average over multiple parameter settings, leads to better performance.
Bayesian deep learning (Neal 2012) provides mechanisms, for building predictive models more robust to adversarial attacks, resilient to unseen or out-of-distribution data and a theoretical framework for estimating model uncertainty (Liu et al. 2019; Ye and Zhu 2018; Wilson et al. 2022; Izmailov et al. 2020; Chen and Ghattas 2020). In particular, quantifying uncertainty enables more reliable decision-making and facilitates identifying potential model vulnerabilities in uncertain regions of the input space. Consequently, embracing Bayesian deep learning (BDL) represents a significant stride towards building more reliable and trustworthy AI systems for various real-world applications (e.g., autonomous driving, medical image analysis, etc.). Unfortunately, their practical use is encumbered by their computational complexity.
Unlike traditional alternatives with point estimates—a single set of model parameters mapping inputs to outputs—Bayesian models learn the distribution of model parameters to offer a distribution over possible predictions. Consider a neural network $f(\mathbf{x},\boldsymbol{\theta})$ with input $\mathbf{x}$ parameterized by $\boldsymbol{\theta}$ and its corresponding prior distribution $p(\boldsymbol{\theta})$. The likelihood $p(\mathcal{D}|\boldsymbol{\theta})$ is determined by $f$ (Bishop 2006) and then, Bayesian inference seeks to use Bayes’ theorem to derive a posterior distribution given by $p\left(\boldsymbol{\theta}|\mathcal{D}\right)={p\left(\mathcal{D}|\boldsymbol{\theta}\right)p\left(\boldsymbol{\theta}\right)}/{p(\mathcal{D})}$ to compute the predictive distribution:
@@
Despite evidence to support the adoption of Bayesian learning, their widespread use is hindered by several challenges; a primary issue is the intractability of the posterior distribution $p(\boldsymbol{\theta}|\mathcal{D})$ in practical learning tasks. The exact solution for the posterior, even for networks of moderate size, is impractical. Because of deep neural network’s complexity and the high-dimensional integral of the resulting denominator. This necessitates the use of approximations.
Our work expands upon recent research demonstrating the effectiveness of ensemble methods and Stein Variational Gradient Descent (SVGD) (Liu and Wang 2016) approaches in BDL (e.g., (Seligmann et al. 2024; Lakshminarayanan, Pritzel, and Blundell 2017; Abbasnejad et al. 2020a; Doan et al. 2022; Izmailov et al. 2020; Chen and Ghattas 2020)) to provide a practical method for ensemble Bayesian deep learning for large-scale tasks. To achieve a practical and efficient approach, we consider spawning parameter particles (models) by adding linear interpolation to a pre-trained model’s parameters to an ensemble. To further enhance the efficiency of these interpolations, inspired by the success of low-rank adapters (LORAs)(Hu et al. 2022), we propose using low-rank parameters in the interpolation process, called Bayesian Low-Rank Learning (Bella).
To validate this, we compare Bella to the full-parameter alternative in terms of (1) accuracy, (2) out-of-distribution generalization, (3) adversarial robustness, and (4) alignment of uncertainty estimation with human confidence. In summary, our findings indicate that this approach achieves comparable performance to full SVGD or ensemble methods at a fraction of the computational cost, with the exception of adversarial robustness. For example, Figure 1 illustrates the application of our proposed approach of Bayesian inference in an application with distribution shifts (CAMELYON17 (Koh et al. 2021) benchmark). The particles (i.e., parameter samples from the posterior) obtained using Bella prove effective in improving generalization, improving the common single particle fine-tuning and comparable with full ensemble-based baselines at a fraction of the cost. Our key contributions are:
@@
  • We propose a new Bayesian learning framework, Bella, for SVGD approximation of a posterior—exploiting the availability of pre-trained models, we spawn particles or models by linear interpolation of a constrained set of model parameters for a SVGD approximation of a posterior.
@@
  • Our approach more efficiently captures the complexity and multi-modality of the solution space compared to current SVGD but at a fraction of the cost—we observe on-par performance with full SVGD in uncertainty estimation, performance improvement, and robustness but only use less than 0.3% of parameters.
@@
  • We demonstrate Bella to perform on par or better than baselines—ensembles or current SVGD implementations. Bella consistently outperforms the non-BNN counterparts on 8 datasets; image classification (ImageNet, CIFAR 10/100), Out-of-distribution (DomainNet, Camleyon17, CIFAR-10-C), and VQA as well as adversarial robustness. Notably, it sets a new state-of-the-art for the CAMELYON17 (see Table 7 for more details) according to the leaderboard (Challenge 2024).
  • Bella employed with the multi-modal model LlaVa (Liu et al. 2023b), leads to improved performance and uncertainty estimation highly correlated with human confidence.
@@ 2 Related Work
Parameter-Efficient Fine Tuning. In contrast to fine-tuning all parameters, recent research proposed inserting adapters in between existing neural layers to reduce the number of trainable parameters and, subsequently, the compute (GPU consumption) (Houlsby et al. 2019; Rebuffi, Bilen, and Vedaldi 2017; Lin, Madotto, and Fung 2020). Hu et al. (Hu et al. 2022) use a bottleneck structure to impose a low-rank constraint on the weight updates, named LoRA. The key functional difference is that LoRA can be merged with the main weights during inference, thus avoiding the introduction of any latency whilst significantly reducing the number of parameters.
@@
Fine-Tuning Approaches and Bayesian Deep Learning.  Previous research investigating the application of fine-tuning approaches for BDLs have predominantly focused on large language models (LLMs), e.g.  (Fan et al. 2020; Zhang et al. 2021; Yang et al. 2024). Notably, marking a departure from the conventional methods relying primarily on tuning the network’s parameters, these studies chose to define priors and approximate posterior over low rank attention weights. Concurrently, Yang et al. (Yang et al. 2024) introduced the concept of Laplace LoRA to incorporate Bayesian concepts to enhance the calibration of fine-tuned LLMs. However, Laplace’s method (Daxberger et al. 2021) relies on a Gaussian approximation of the posterior distribution. Whilst this can be effective for unimodal and symmetric distributions, the approach does not fully encapsulate the intricacies of more complex posteriors, particularly in neural networks where the posterior has multimodality and asymmetry (Izmailov et al. 2021). Deep ensembles (Lakshminarayanan, Pritzel, and Blundell 2017) typically perform better in practice compared with variational and Laplace methods, due to their ability to capture multiple modes. When employing fine-tuning, a direct application of ensembling for LoRAs was considered in (Wang, Aitchison, and Rudolph 2024). Some interpretations of deep ensembles suggest that they approximate gradient flows in function spaces and that building desirable properties into an ensemble (such as repulsive behavior), is possible (Wild et al. 2023). SVGD (Liu and Wang 2016), can be viewed in a similar vein. However, while these more sophisticated, repulsive, ensembling approaches are highly impractical in the pre-training phase, we argue that their expressivity can be brought to bear precisely in tandem with low-rank fine-tuning, which is the viewpoint we adopt in this contribution.
@@
Building upon these foundations, our research represents a pioneering effort to apply the principles of repulsive ensemble-based low-rank fine-tuning to computer vision. In particular, we bridge pre-training and fine-tuning phases with recent conjectures on mode connectivity (Ainsworth, Hayase, and Srinivasa 2023). Our methodology not only capitalizes on the efficiency of fine-tuning techniques, e.g. (Ding et al. 2023, 2022; Dettmers et al. 2023) but also innovatively addresses the scalability challenges inherent to BDLs. Overall, our work sets a new precedent in applying Bayesian approaches to computer vision tasks by offering a scalable and efficient framework for enhancing model performance.
3 Background on Stein Variational Gradient Descent (SVGD)
@@
Bayesian inference techniques have been integral to the development of neural networks, with a rich history underscored by previous works (Neal et al. 2011; Neal 2012; MacKay 1991; Welling and Teh 2011; Blei, Kucukelbir, and McAuliffe 2017). Variational Inference (VI) (Blundell et al. 2015; Blei, Kucukelbir, and McAuliffe 2017) and Markov Chain Monte Carlo (MCMC) (Neal et al. 2011; Welling and Teh 2011) are two primary approximate Bayesian inference frameworks. The former substitutes the true posterior with a tractable alternative while the latter involves sampling. However, accurately computing the posterior with either MCMC or VI becomes computationally infeasible when dealing with large-scale networks containing millions of parameters. Although approximations can be obtained more efficiently with VI (Blundell et al. 2015), VI is also demonstrably too restrictive to resemble the multi-modality of the true posterior and suffers from mode collapse (Izmailov et al. 2021).
@@
Stein Variational Gradient Descent (SVGD) (Liu and Wang 2016), among other ensemble-based approaches, is an alternative approximate Bayesian technique that combines the strengths of MCMC and VI by transporting a set of parameter particles to fit the true posterior distribution, while encouraging diversity among the particles, by incorporating a repulsive term in the parameter updates. This diversity prevents the mode collapse and enables learning multiple models to represent various patterns in the data. Using $n$ samples from the posterior (i.e. parameter particles), SVGD modifies the gradient descent as:
$\boldsymbol{\theta}_{i}$ $=\boldsymbol{\theta}_{i}-\epsilon_{i}\hat{\boldsymbol{\phi}}{}^{*}(\boldsymbol{\theta}_{i})\text{\quad with}$
$\hat{\boldsymbol{\phi}}{}^{*}(\boldsymbol{\theta})$ $=\sum_{j=1}^{n}\big{[}k(\boldsymbol{\theta}_{j},\boldsymbol{\theta})\nabla_{\boldsymbol{\theta}_{j}}\log p(\boldsymbol{\theta}_{j}|\mathcal{D})-\frac{\gamma}{n}\nabla_{\boldsymbol{\theta}_{j}}k(\boldsymbol{\theta}_{j},\boldsymbol{\theta})\big{]}\,.$
@@
4 Bayesian Low-Rank Learning (Bella)
The problem with the current SVGD and other ensemble-based methods in large deep neural networks is its huge computational cost. This renders it infeasible to train efficiently and to scale to a sufficient number of parameter particles for accurately approximating the posterior distribution, which currently remains coarse. In this work, we propose to capitalize on the low-rank representations of fine-tuning in order to construct a practical and scalable variant of SVGD. We note that while our approach may not fully capture the diversity of the multi-modal posterior, a recent conjecture (Ainsworth, Hayase, and Srinivasa 2023) suggests that these modes might result from parameter permutations in neural networks, leading to models that are functionally equivalent. Building on this idea, Bella could serve as a practical alternative with sufficient theoretical justification for ensemble methods.
Consider any dense layer, for which there is a fixed pre-trained weight matrix $\boldsymbol{\theta}_{0}\in\mathbb{R}^{d_{1}\times d_{2}}$ with $d_{1},d_{2}$ the corresponding numbers of hidden units. We consider $n$ low-rank perturbations of $\boldsymbol{\theta}_{0}$ as